# Using Dynamic Safety Margins as Control Barrier Functions

### v2

@@ 6 Examples
In this section, we present two nonlinear examples to showcase the presented approach of designing CBFs using dynamic safety margins. For each example, we provide a comparison with candidate CBFs and the ERG. Throughout this section, simulations are computed in MATLAB and optimization problems are solved with MOSEK [34].
@@ 6.1 Anthill system
$\dot{x}=(x^{2}-1)x+u,$ (41)
with state space domains $\mathbb{X}=\mathbb{R}$ and $\mathbb{U}=\mathbb{R}$. In the absence of inputs $u=0$, this system asymptotically tends to 00 if $|x_{0}|<1$ and diverges to infinity if $|x_{0}|>1$. Let $\mathcal{U}=[-u_{\max},u_{\max}]$ be the input constraint set, with $u_{\max}\!=\!\frac{4}{3\sqrt{3}}$.
@@
with $v\in\mathbb{V}=\mathbb{R}$, the prestabilized system is
@@
which satisfies the equilibrium conditions $\bar{x}(v)=v$ and $\bar{u}(v)=\pi\big{(}\bar{x}(v),v\big{)}=v-v^{3}$ for all $v\in\mathbb{V}$. The equilibrium point $\bar{x}(v)$ is attractive if $|x_{0}-v|<1$. This can be proven using the Lyapunov function
@@
This anthill-shaped function (see Fig. 2) is such that, $\forall v\in\mathbb{V}$,
$V\big{(}\bar{x}(v),v\big{)}$ $=0,$
$V(x,v)$ $>0,\quad\forall x\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(v)\},$
$\frac{\partial V}{\partial x}f_{\pi}(x,v)$ $=-f_{\pi}^{2}(x,v)<0,\quad\forall x\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(v)\},$
where $\mathcal{D}_{\boldsymbol{\mathbf{v}}}=(v-1,v+1)$ is the basin of attraction of the equilibrium point $\bar{x}(v)$. The function $V(x,v)$ also satisfies Assumption 4. The stability threshold value (36) is
@@
To obtain the steady-state admissible set $\mathcal{V}\subset\mathbb{V}$, consider the input constraints at steady-state $|\bar{u}(v)|=|v-v^{3}|\leq u_{\max}$. Using Cardano’s formula to compute the roots of a 3-rd degree polynomial, we obtain $\mathcal{V}=\left[-v_{\max},v_{\max}\right]$, where
@@
The resulting threshold value is
@@
Noting that $V\big{(}\frac{1}{2}v+\delta^{+}(v),v\big{)}=\frac{1}{4}=\overline{\Gamma}(v)$ if $v=v_{\max}\pm 1$, it can be shown that
$\Gamma(v)=0.99\max\left(\min\left(\Gamma^{*}(v),\overline{\Gamma}(v)\right),-\overline{\Gamma}(v)\right),$ (48)
is a continuously differentiable function satisfying the conditions of Theorem 5.26. Thus, $\Delta(x,v)=\Gamma(v)-V(x,v)$ is a Lyapunov-based DSM for $\pi$ and a CBF by Theorem 3.7. Furthermore, the set $\mathcal{C}\subset\mathbb{X}\times\mathcal{V}$ is compact in $\mathbb{X}\times\mathbb{V}$ (see Remark 5.28).
Figure 2: Anthill-shaped Lyapunov function $V(x,v)$ at $v=0.5$ and the associated threshold value $\Gamma^{*}(v)$ and stability threshold value $\overline{\Gamma}(v)$.
We now consider the state constraints $\mathcal{X}=[-x_{\max},x_{\max}]$, with $x_{\max}=2$. Since $x_{\max}>v_{\max}$ and (41) is a first order system (no overshoot), these constraints have already been enforced by the restriction $v\in\mathcal{V}$ induced by $\mathcal{C}$. Thus, no further action is needed.
@@
    DSM-based CBF: We use $\eta=0.01$, the class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 1.8c$, and $\rho:(v,r)\mapsto r-v$.
@@
The simulation results for each approach are shown in Fig. 3. Both the ERG and the proposed DSM-CBF successfully enforce safety and reach the closest safe point to the desired reference but, the DSM-CBF exhibits performance that more closely resembles the nominal behavior. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. Conversely, the DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the stability region of the nominal or prestabilizing controllers.
Figure 3: Closed-loop behavior for the anthill system using each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and the DSM-CBF approaches, respectively. Note that $v(t)$ may decrease in our approach, which is not possible in the ERG. This feature enlarges the feasible region of the CBF-based QP and allows the safe input $\beta(x,v)$ to be closer to the nominal input $\pi(x,r)$. In addition, we show the performance of the candidate CBFs for different class $\mathscr{K}_{\infty}$ functions $\alpha_{1}=\alpha_{2}=\alpha:c\mapsto a_{i}c$, where $a_{i}\in\{7,0.25,0.15,0.1,0.07\}$. Larger $a_{i}$ corresponds to better input tracking, but also an earlier failure due to infeasibility, marked by a magenta ‘$\times$’ marker.
@@ 6.2 Overhead Crane
where the degrees of freedom $\boldsymbol{\mathbf{q}}=[x;~{}\theta]$ are the gantry position $x$ and the payload angle $\theta$, and
@@
[…]
@@
    $\Gamma_{1}(v)$ $=\Gamma_{1}^{*}(v)=\frac{1}{2}k_{p}(v-x_{\text{min}})|v-x_{\text{min}}|,$
    $\Gamma_{2}(v)$ $=\Gamma_{2}^{*}(v)=\frac{1}{2}k_{p}(x_{\text{max}}-v)|x_{\text{max}}-v|.$
@@
    $\Gamma_{3}(v)=\Gamma_{3}^{*}(v)=\frac{m_{c}u_{\text{max}}^{2}}{2(m_{c}k_{p}+k_{d}^{2})}.$
@@
    $\Gamma_{4}(v)=\Gamma_{4}^{*}(v)=m_{p}gL(1-\cos\theta_{\text{max}}).$
    We note that this threshold value is strictly smaller than the stability threshold value (36), i.e., $\overline{\Gamma}(v)=2m_{p}gL$. As such, enforcing the angle constraints ensures $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{v}$.
@@
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [28, Proposition 4], we can define
@@
    which holds $\forall x\leq x_{\text{max}}$ and $\forall\theta\in[-\pi/2,\pi/2]$. With these approximations, the following function satisfies the condition of Theorem 5.26:
@@
The resulting DSMs are $\Delta_{i}(\boldsymbol{\mathbf{x}},v)=\Gamma_{i}(v)-V(\boldsymbol{\mathbf{x}},v)$ for all $i\in\{1,\ldots,5\}$. By Theorem 4.22, these DSMs have the control-sharing property. For all simulations, we use linear class $\mathscr{K}_{\infty}$ functions given by $\alpha_{i}:c\mapsto\alpha_{i}c$ where $\alpha_{i}>0$ represents both the function and the scalar gain despite the slight abuse of notation. We choose $\alpha_{1}=\alpha_{2}=10$, $\alpha_{3}=20$, $\alpha_{4}=80$ and $\alpha_{5}=2.2$. The navigation field we use is $\rho(v,r)=r-v$.
The following approaches are used for comparison.
@@
    $\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\kappa(\boldsymbol{\mathbf{x}},v)\\ \min_{i\in\{1,\ldots,5\}}\Delta_{i}(\boldsymbol{\mathbf{x}},v)\rho_{\text{ERG}}(v,r)\end{bmatrix}.$ (52)
@@
    Where $\gamma_{i}$ are positive constants. For all simulations, we use $\gamma_{i}=6$ and $\tilde{\alpha}_{i}=8,\forall i\in\{1,\ldots,4\}$, and $\gamma_{5}=4$, $\tilde{\alpha}_{5}=3.5$, where $\tilde{\alpha}_{i}:c\mapsto\tilde{\alpha}_{i}c$ is the class $\mathcal{K}_{\infty}$ function associated with $h_{i}$. It must be noted that all $h_{i}$ described above are candidate CBFs, meaning that they may not satisfy Definition 2 or have the control-sharing property given in Definition 4.21.
@@
$\eta$ (-) 0.01 $k_{d,\kappa}$ $4$ $\theta_{\text{max}}$ (deg) $10$ $r$ (m) $1$
General parameters for the example are provided in Table 1. For the first simulation, shown in Fig. 5, we relax the angle constraint to $|\theta|\leq\pi/2$, which effectively makes it inactive. In this case, all constrained control approaches achieve the safety objective. In terms of performance, the Candidate CBF approach tracks the nominal behavior better than the proposed DSM-CBF approach, whereas the ERG exhibits the slowest response. The second simulation, shown in Fig. 6, includes the angle constraint $|\theta|\leq\theta_{\text{max}}$. The ERG and our proposed DSM-CBF approaches successfully enforce the constraints, with the proposed method achieving better performance. However, the Candidate CBF-based controller becomes infeasible despite our best efforts to tune $\gamma_{i}$ and $\tilde{\alpha}_{i}$. This is unsurprising since candidate CBFs do not guarantee persistent feasibility.
Figure 5: Closed-loop behavior in the absence of angle constraints. Even though the nominal controller violates the constraints, the Candidate CBF, ERG, and proposed DSM-CBF approaches all successfully enforce the safety requirements.
Figure 6: Closed-loop behavior in the presence of angle constraints. The ‘$\times$’ marks the instant in which the Candidate CBF policy becomes infeasible, likely due to the unavoidable backswing experienced by the payload angle $\theta(t)$. Conversely, the ERG and proposed DSM-CBF approaches successfully enforce the safety requirements.
@@ 7 Conclusion
This paper showed how dynamic safety margins can be used to systematically construct control barrier functions with the control-sharing property. The work also highlighted a Lyapunov-based approach to compute DSM-CBFs and presented simulation results that demonstrate their ability to guarantee that the CBF optimization problem is persistently feasible for systems with high relative degree, multiple state and input constraints, and that are vulnerable to a loss of control authority. Future work includes continuing to explore connections between the CBF and reference governor literatures and robustifying the presented results to account for bounded disturbances.

### v3

@@ 6 Examples
In this section, we present two nonlinear examples to showcase the presented approach of designing CBFs using dynamic safety margins. For each example, we provide a comparison with candidate CBFs, backup CBFs, and the ERG. Simulations are computed in MATLAB and optimization problems are solved with MOSEK [40].
@@ 6.1 Anthill system
$\dot{x}=(x^{2}-1)x+u.$ (40)
In the absence of inputs $u\!=\!0$, the system asymptotically tends to 00 if $|x_{0}|\!<\!1$ and diverges to infinity if $|x_{0}|>1$. Let $\mathcal{U}=[-u_{\max},u_{\max}]$ be the input constraint set, with $u_{\max}\!=\!\frac{4}{3\sqrt{3}}$.
@@
with $v\in\mathbb{R}$, the prestabilized system is
@@
which satisfies the equilibrium conditions $\bar{x}(v)=v$ and $\bar{u}(v)=\pi\big{(}\bar{x}(v),v\big{)}=v-v^{3}$ for all $v\in\mathbb{R}$. The equilibrium point $\bar{x}(v)$ is attractive if $|x_{0}-v|<1$. This can be proven using the Lyapunov function
@@
This anthill-shaped function (see Fig. 2) is such that, given the $\mathscr{K}_{\infty}$ functions $\gamma_{1}(c)=\frac{1}{4}c^{2}$ and $\gamma_{2}(c)=\frac{1}{2}c^{2}$, $\forall(x,v)\in\tilde{\mathcal{D}}=\{(x,v)\in\mathbb{R}\times\mathbb{R}:x\in(v-1,v+1)\}$,
$\gamma_{1}(|x-v|)$ $\leq V(x,v)\leq\gamma_{2}(|x-v|),$
$\frac{\partial V}{\partial x}f_{\pi}(x,v)$ $=-f_{\pi}^{2}(x,v)\leq 0,$
The function $V(x,v)$ also satisfies Assumption 4. The stability threshold value (36) is
@@
To obtain the steady-state admissible set $\mathcal{V}\subset\mathbb{R}$, consider the input constraints at steady-state $|\bar{u}(v)|=|v-v^{3}|\leq u_{\max}$. Using Cardano’s formula to compute the roots of a 3-rd degree polynomial, we obtain $\mathcal{V}=\left[-v_{\max},v_{\max}\right]$, where
@@
The resulting safety threshold value is
@@
While this function is discontinuous, it is bounded by the smooth function $\Gamma_{s}^{*}(v)=\max\left(\min\big{(}\Gamma^{*}(v),\frac{1}{4}\big{)},-\frac{1}{4}\right)$. Thus, by Theorem 4.25, $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{2}$ defined as
$\Delta(x,v)=\begin{bmatrix}\Gamma_{s}^{*}(v)-V(x,v)\\ 0.99\overline{\Gamma}(v)-V(x,v)\end{bmatrix},$ (47)
is a DSM for $\pi$ and, by Theorem 3.6, a (vector-valued) CBF with respect to the augmented system.
Figure 2: Anthill-shaped Lyapunov function $V(x,v)$ at $v=0.5$, with the safety threshold value $\Gamma^{*}(v)$ and stability threshold value $\overline{\Gamma}(v)$.
We now consider the state constraints $\mathcal{X}=[-x_{\max},x_{\max}]$, with $x_{\max}=2$. Since $x_{\max}>v_{\max}$ and (40) is a first order system (no overshoot), these constraints have already been enforced by the restriction $v\in\mathcal{V}$ induced by $\tilde{\mathcal{C}}$. Thus, no further action is needed.
@@
    DSM-CBF: We use $\eta=0.01$, the class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 1.8c$, and $\rho:(v,r)\mapsto r-v$.
@@
  4. (d)
    Backup CBF: We consider the backup policy
    $\psi(x)=u_{\max}\left(\frac{1-e^{10x}}{1+e^{10x}}\right),$ (48)
    and formulate the Backup CBF
    $h(x)=\min_{\tau\in[0,10]}x_{\max}-\Phi_{f_{\psi}}(x,\tau),$ (49)
    where $\Phi_{f_{\psi}}(x_{0},\tau)$ is the solution of the backup dynamics $\dot{x}=f_{\psi}(x)\triangleq(x^{2}-1)x+\psi(x)$ at time $t=\tau$ with initial condition $x(0)=x_{0}$. The implementation is done following [21] with class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 10c$.
The simulation results for each approach are shown in Fig. 3. Both the ERG and the proposed DSM-CBF successfully enforce safety and reach the closest safe point to the desired reference but, the DSM-CBF exhibits performance that more closely resembles the nominal behavior. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. The backup CBF also fails in this example. Although the backup policy $\psi$ is safe for the starting condition considered, it loses control authority as soon as $x>v_{\max}$, which causes it to become infeasible. Conversely, the DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the input saturation.
Figure 3: Closed-loop behavior for the anthill system using each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and the DSM-CBF approaches, respectively. In addition, we show the performance of the candidate CBFs for different class $\mathscr{K}_{\infty}$ functions $\alpha_{1}=\alpha_{2}=\alpha:c\mapsto a_{i}c$, where $a_{i}\in\{7,0.25,0.15,0.1,0.07\}$. Larger $a_{i}$ corresponds to better input tracking, but also an earlier failure due to infeasibility, marked by a magenta ‘$\times$’ marker. The backup CBF is shown in orange and the ‘$\times$’ marker signals the point when the backup policy becomes unsafe, invalidating the approach.
@@ 6.2 Overhead Crane
where the generalized coordinates $\boldsymbol{\mathbf{q}}=[x;~{}\theta]$ are the gantry position $x$ and the payload angle $\theta$, and
@@
which satisfies Definition 4.21 with $\mathcal{D}_{v}=\mathbb{R}\times(-\pi,\pi)\times\mathbb{R}^{2}$ and also satisfies Assumption 4. We now define four types of constraints and, for each one, we find the associated safety threshold value or a continuously differentiable bounding function. For details on the construction of each $\Gamma^{*}_{i}(v)$, the reader is referred to [30, 39].
@@
    $\Gamma_{1}^{*}(v)$ $=\frac{1}{2}k_{p}(v-x_{\text{min}})|v-x_{\text{min}}|,$
    $\Gamma_{2}^{*}(v)$ $=\frac{1}{2}k_{p}(x_{\text{max}}-v)|x_{\text{max}}-v|.$
@@
    $\Gamma^{*}_{3}(v)=\frac{m_{c}u_{\text{max}}^{2}}{2(m_{c}k_{p}+k_{d}^{2})}.$
@@
    $\Gamma_{4}^{*}(v)=m_{p}gL(1-\cos\theta_{\text{max}}).$
@@
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [30, Proposition 4], we can define
@@
    which holds $\forall x\leq x_{\text{max}}$ and $\forall\theta\in[-\pi/2,\pi/2]$. With these approximations, the following continuously differentiable function is a bound for $\Gamma^{*}_{5}(v)$ and can be used in its stead:
@@
The stability threshold value is
$\overline{\Gamma}(v)=2m_{p}gL.$
Using Theorem 4.25, we conclude that
$\Delta(\boldsymbol{\mathbf{x}},v)=\begin{bmatrix}\Gamma^{*}_{1}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{2}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{3}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{4}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma_{5}(v)-V(\boldsymbol{\mathbf{x}},v)\\ 0.99\overline{\Gamma}(v)-V(\boldsymbol{\mathbf{x}},v)\end{bmatrix},$
is a DSM for $\pi$ and, by Theorem 3.6, a (vector-valued) CBF with respect to the augmented system. Note that the stability threshold value satisfies $\Gamma^{*}_{4}(v)<\overline{\Gamma}(v)$. As such, when the angle constraints are enforced, there is no need to include $\overline{\Gamma}(v)$ in the definition of $\Delta$ above.
For all simulations, we use linear class $\mathscr{K}_{\infty}$ functions given by $\alpha_{i}:c\mapsto\alpha_{i}c$ where $\alpha_{i}>0$ represents both the function and the scalar gain despite the slight abuse of notation. We choose $\alpha_{1}=\alpha_{2}=10$, $\alpha_{3}=20$, $\alpha_{4}=80$, $\alpha_{5}=2.2$ and $\alpha_{6}=80$. The navigation field we use is $\rho(v,r)=r-v$.
We compare with the following approaches.
@@
    $\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\kappa(\boldsymbol{\mathbf{x}},v)\\ \min_{i}\Delta_{i}(\boldsymbol{\mathbf{x}},v)\rho_{\text{ERG}}(v,r)\end{bmatrix}.$ (53)
@@
    Where $\gamma_{i}$ are positive constants. For all simulations, we use $\gamma_{i}=6$ and $\tilde{\alpha}_{i}=8,\forall i\in\{1,\ldots,4\}$, and $\gamma_{5}=4$, $\tilde{\alpha}_{5}=3$, where $\tilde{\alpha}_{i}:c\mapsto\tilde{\alpha}_{i}c$ is the class $\mathcal{K}_{\infty}$ function associated with $h_{i}$. It must be noted that all $h_{i}$ described above are candidate CBFs, meaning that they may not satisfy Definition 2 or have the control-sharing property given in Definition 2.4.
  4. (d)
    Backup CBF: We consider the backup policy
    $\psi(\boldsymbol{\mathbf{x}})=-u_{\max}.$ (54)
    For each constraint, we design a backup CBF and implement it according to [21]. For instance, the backup CBF associated to the payload constraint is
    $h(\boldsymbol{\mathbf{x}})\!=\!\min_{\tau\in[0,10]}p_{\max}\!-\!\boldsymbol{\mathbf{e}}_{1}^{\top}\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\!-\!L\sin\left(\boldsymbol{\mathbf{e}}_{2}^{\top}\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\right),$
    where $\boldsymbol{\mathbf{e}}_{i}$ is the $i$-th column of $I_{4}$ and $\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}}_{0},\tau)$ is the solution of the backup dynamics $f_{\psi}$ at time $t=\tau$ with initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$. All class $\mathscr{K}_{\infty}$ functions are the same $\alpha(c)=300c$.
@@
$\eta$ (-) 0.02 $k_{d,\kappa}$ $4$ $\theta_{\text{max}}$ (deg) $10$ $r$ (m) $1$
General parameters for the example are provided in Table 1. For the first simulation, shown in Fig. 5, we remove the angle constraint. In this case, all constrained control approaches achieve the safety objective. In terms of performance, the backup and candidate CBF approaches track the nominal behavior better than the proposed DSM-CBF approach, whereas the ERG exhibits the slowest response. The second simulation, shown in Fig. 6, includes the angle constraint $|\theta|\leq\theta_{\text{max}}$. The ERG and our proposed DSM-CBF approaches successfully enforce the constraints, with the proposed method achieving better performance. However, the candidate CBF-based controller becomes infeasible despite our best efforts to tune $\gamma_{i}$ and $\tilde{\alpha}_{i}$. This is unsurprising since candidate CBFs do not guarantee persistent feasibility. The backup policy $\psi$ is unsafe under the introduced angle constraints. This highlights that finding a suitable backup policy becomes more challenging as the number of constraints increases.
Figure 5: Closed-loop behavior in the absence of angle constraints. Even though the nominal controller violates the constraints, the ERG, candidate CBF, backup CBF and proposed DSM-CBF approaches all successfully enforce the safety requirements. The dashed blue and green lines are the virtual reference $v$ of the DSM-CBF and ERG approaches, respectively. Note that $v(t)$ may decrease in our approach, which is not possible in the ERG. This feature enlarges the feasible region of the CBF-based QP and allows the safe input $\beta(x,v)$ to be closer to the nominal input $\pi(x,r)$.
Figure 6: Closed-loop behavior in the presence of angle constraints. The ‘$\times$’ marks the instant in which the candidate and backup CBF policies become infeasible. The failure in the candidate CBF is likely due to the unavoidable backswing experienced by the payload angle $\theta(t)$. The failure in the backup CBF is due to the backup policy being unsafe with respect to the angle constraints. Conversely, the ERG and proposed DSM-CBF approaches successfully enforce the safety requirements.
@@ 7 Conclusion
This paper showed that dynamic safety margins are control barrier functions for an augmented system obtained by concatenating the state and reference vectors. The work presents an optimization-based safe policy and it is shown that it is persistently feasible. Then, a Lyapunov-based approach to construct DSMs was presented in detail and its performance against other common approaches to design CBFs, as well as against the ERG was compared in two nonlinear systems.
Future work includes continuing to explore connections between the CBF and reference governor literatures, adapt other types of DSMs to construct CBFs, and robustifying the presented results to account for bounded disturbances.